What is the answer evaluate the function y=3(1/2)* for x=-3.
1
Expert's answer
2020-05-19T18:46:27-0400

"y(x)=3(\frac{1}{2})^x"

"y(-3)=3(\frac{1}{2})^{-3}=3*2^3=24."